# Differentiable Optimization of Similarity Scores Between Models and Brains

### v1

@@ 4.2.2 Neural datasets
The optimization dynamics revealed in Figure 4a for a two-dimensional dataset also holds on real neural datasets (A.1). We now consider various neural datasets as the reference dataset for the optimization procedure. Figure 5a shows the optimization dynamics when a random noise Gaussian dataset is optimized towards the Siegel et al. 2015 dataset by maximizing angular Procrustes similarity (top) or CKA similarity (bottom). Each curve shows how the optimized dataset captures a single principal component of the reference dataset during the course of optimization. Similar to the results in Figure 4, optimizing for angular Procrustes similarity captures more of the lower variance components in the data for a given similarity score.
@@
A convenient way of summarizing these curves is to note the similarity score required to capture a given principal component above some threshold, defined here (and shown in Figure 5a) as the centerpoint between the maximum and initial $R^{2}$ value for a given principal component. Figure 5b shows the score required to reach the principal component threshold for the different principal components in the Mante et al. 2013 dataset. Figures 5c and 5d show the same curves when the reference datasets are now the electrode recordings from Siegel et al. 2015 and Majaj et al. 2015.
Angular CKA. Williams et al. (2021) proposed to take the arccosine of CKA to obtain a measure that satisfies the axioms of a distance metric. Figures 5b, 5c, and 5d additionally show that the metric version of CKA (referred to as angular CKA) improves its sensitivity to lower variance components.
@@
Neural curves predicted by matched Gaussians. Surprisingly, the optimization dynamics shown in these figures is well matched when the reference neural dataset is replaced by a new reference dataset that consists of random Gaussian numbers with variances for each principal component matched to the neural data (Figure 5e). The variance distribution of the reference dataset strongly determines when each principal component is captured during optimization.
4.3 Theoretical analysis of CKA and NBS
Ding et al. (2021) briefly noted in their discussions that in the simplified case where two datasets have the same eigenvectors, CKA reduces to products of squared singular values of each dataset, whereas Procrustes distance doesn’t square the singular values. Here we provide a more formal way to derive this result, leveraging a duality relationship discovered by Harvey et al. (2023).
Duality of Bures and shape distances. Harvey et al. (2023) showed that angular Procrustes is related to another measure, Normalized Bures Similarity (NBS) (Tang et al., 2020), by just taking the arccosine. Similar to angular CKA, we find that the arccosine increases sensitivity to lower variance components (Figures 5b, 5c, and 5d). This can be understood by the nonlinear mapping of arccosine that expands values close to one.
To understand the increased sensitivity to low variance principal components of NBS compared to CKA (dark blue and red curves in Figures 5b, 5c, and 5d), we analyze their matrix norm formulation. As shown in Harvey et al. (2023), CKA and NBS can be formulated as follow (see appendix B.1 for details):
@@ 4.3 Theoretical analysis of CKA and NBS
NBS quantifies similarity using the nuclear norm, which involves a sum of singular values, whereas CKA uses the Frobenious norm, which sums the square of the singular values. The additional square operation in CKA significantly increases the contribution of large variance components. We show this in theory by considering how CKA and NBS change when perturbing a single principal component of the data. The result is that CKA depends quadratically on the variance of the perturbed principal component, whereas NBS has a linear dependence (see proof in appendix B.3), as empirically validated in Figure 6a.
@@
where $\tilde{X}_{k}$ is equal to $X$ where only the $k^{\text{th}}$ principal component is modified.
@@ 4.4 Are metrics mutually independent?
A natural question that arises is whether these different similarity metrics are independent of each other, or if they exhibit consistent relationships. To address this, we consider three possible relationships between any two given similarity metrics (Figure 6b):
  • Independent: The two metrics are independent. A high score on one metric offers no guarantee of a high score on the other.
  • Coupled: The two metrics are tightly coupled. A high score on one metric implies a high score on the other, and vice versa.
  • Asymmetric: One metric subsumes the other. A high score on the first metric guarantees a high score on the second, but not the other way around.
We jointly optimize multiple similarity scores to find their allowed ranges (Figure 6c) and show that a high angular Procrustes similarity implies a high CKA score, but not the converse. A high value of angular Procrustes implies a high score for unregularized linear regression but linear regression that is regularized and cross-validated across experimental conditions can take independent values.
Figure 6: (a) The similarity score between a synthetic dataset and a modified version of the same dataset when a single principal component is perturbed. For all similarity measures, when small variance components are perturbed the similarity score is near 1. However, when high variance components are perturbed the angular Procrustes score drops much more than for CKA and Normalized Bures Similarity (NBS). (b, c) We jointly optimized the values of both angular Procrustes and CKA or linear regression to illustrate the allowed ranges of both similarity scores (gray region enclosed by the solid lines). The pink dots in panel c indicate the target values for optimization, and the blue curves indicate the optimization trajectories of the two similarity measures. If angular Procrustes has a high score of 0.9 (horizontal dashed line) then linear regression will have a value above this. In contrast, a high linear regression score of 0.9 (vertical dashed line) does not imply a high angular Procrustes score, and a wide range of angular Procrustes scores are possible.
@@ 5 Discussion
Limitations & future work. Our study focused on differentiable, geometry-based similarity measures and their optimization dynamics. Future work should investigate how it extends to other types of measures. Additionally, a comprehensive analysis of computational efficiency and scalability for different metrics will be crucial as datasets continue to grow in size. Moving forward, our python package aims to standardize similarity measures, facilitating a more cumulative scientific approach by centralizing findings related to these measures, and enabling more integrated benchmarking and comparisons.
Appendix A Neural Datasets and Models
A.1 Datasets
@@
  • Mante et al. [2013]: Prefrontal cortex (PFC) electrode recordings during a contextual decision-making task involving colored moving dots. Data link111 https://www.ini.uzh.ch/en/research/groups/mante/data.html.
  • Siegel et al. [2015]: Frontal Eye Field (FEF) electrode recordings during a contextual decision-making task involving colored moving dots, similar to Mante et al. (2013).
  • Hatsopoulos et al. [2007]: Primary motor (M1) electrode recordings during a center-out reaching task. Data link [^2] https://datadryad.org/stash/dataset/doi:10.5061/dryad.xsj3tx9cm.
  • Majaj et al. [2015]: Inferior temporal (IT) electrode recordings during object image presentations.
  • Freeman et al. [2013]: Primary (V1) and secondary (V2) visual area electrode recordings during texture and noise image presentations.
We use the BrainScore [^3] https://github.com/brain-score/vision library [Schrimpf et al., 2020] for the Majaj et al. [2015], Freeman et al. [2013] datasets.
A.2 Recurrent Neural Network (RNN) models
We show results for three commonly used RNN architectures: LSTMs [Hochreiter and Schmidhuber, 1997] and two choices for continuous time recurrent neural networks (CTRNNs), which differ by the position of the nonlinearity [Miller and Fumarola, 2012]. A first alternative is given by the following equations.
@@ A.2 Recurrent Neural Network (RNN) models
We call it the LowPassCTRNN since it can be shown that its firing rate is a low-pass filter of the firing rate of the first architecture, which we call CTRNN [Miller and Fumarola, 2012]. The RNNs are trained using supervised learning on simplified task inputs and outputs.
A.3 Condition average baseline:
@@
The condition average baseline is obtained by averaging the neural activity along the condition dimension, while still allowing each neuron to have its own unique time-varying firing rate, and then comparing this condition-averaged dataset to the original neural data. This baseline shows the similarity one can obtain by discarding the condition specific activity to keep only the condition independent dynamics of the neural activity.
Appendix B Similarity Measures
B.1 Definitions
@@
Several methods have been proposed to measure similarity between models and neural data (see [Sucholutsky et al., 2023], [Klabunde et al., 2023] for comprehensive reviews). While some efforts have been made to characterize these metrics mathematically, for instance, by examining their invariance properties, clear guidance on selecting the most appropriate metric for a given scenario remains limited. To address this gap, we introduce a novel method for analyzing the specific aspects of data prioritized by different similarity metrics. Our method leverages the differentiability of these metrics and is applicable to a wide range of measures. We focus here on three commonly used metrics and their variants, linear regression [Yamins et al., 2014, Schrimpf et al., 2018], Centered Kernel Alignment (CKA) [Kornblith et al., 2019], and Procrustes distance [Williams et al., 2021, Ding et al., 2021]. These methods quantify similarity based on the goodness of fit after aligning the representations. This alignment transformation allows for greater flexibility by making similarity measures invariant to specific transformations. For example, instead of demanding a one-to-one mapping of neurons between datasets, these methods can accommodate scenarios where neurons in one dataset correspond to linear combinations of neurons in the other.
@@
Unlike linear regression, CKA and Procrustes are symmetric, meaning the metric yields the same result whether applied to X, Y or Y, X. Moreover, they exhibit a different class of invariance. While linear regression is invariant under invertible linear transformations, CKA and Procrustes are invariant under orthogonal linear transformations. This stricter invariance to orthogonal transformations potentially reduces sensitivity to noise [Kornblith et al., 2019].
CKA measures the correlation between the kernels of two datasets $X$ and $Y$ [Kornblith et al., 2019]. We consider here linear CKA where the kernels are $XX^{T}$ and $YY^{T}$.
@@
CKA scores range from 0 to 1, with 1 indicating perfect similarity. We also consider the arccosine of CKA, a metric satisfying the axioms of a distance metric (e.g., the triangle inequality) with 0 signifying perfect similarity [Williams et al., 2021]. This metric, also known as Angular CKA [Lange et al., 2022], is then normalized to a similarity score between 0 and 1 for direct comparison with other scoring methods. This normalization involves dividing the angular distance by $\pi/2$ and subtracting the result from 1, ensuring the measure increases with similarity.
@@
Procrustes distance provides another approach for quantifying similarity [Williams et al., 2021, Ding et al., 2021]. This metric identifies the optimal orthogonal alignment to maximize the correlation between $X$ and $Y$:
@@
where $\mathcal{O}$ is the group of orthogonal linear transformations. An angular distance metric can be obtained by taking the arccosine of this quantity.
As shown in Harvey et al. [2023], Procrustes distance is closely related to Normalized Bures Similarity (NBS) [Tang et al., 2020], with Procrustes distance equating the arccosine of NBS. NBS is defined as:
@@
This definition resembles CKA but utilizes the nuclear matrix norm instead of the Frobenius matrix norm. The Frobenius norm, denoted by $\|A\|_{F}$ for a matrix $A$, is calculated as the square root of the sum of squared singular values. The nuclear norm, $\|A\|_{*}$, is simply the sum of singular values. The implications of this difference in matrix norms for the weighting of principal components by CKA and NBS are further explored in B.3.
@@
B.2 Optimizing similarity scores
To better characterize similarity measures we optimize synthetic datasets $Y$ to become more similar to a reference dataset $X$. We initialize the synthetic dataset $Y$ by randomly sampling from a standard Gaussian distribution with the same shape as $X$. We use Adam [Kingma and Ba, 2017] to optimize $Y$ to maximize the similarity score with $X$, leveraging the differentiability of the similarity measures, and stop the optimization when the score reaches a fixed threshold.
As we optimize $Y$ towards greater similarity with $X$, we simultaneously evaluate how well task-relevant variables can be linearly decoded from the evolving synthetic data. This decoding analysis employs logistic regression with stratified 5-folds cross-validation. For dataset with temporal dynamics, we fit a separate decoder for each time step and report the average accuracy across time.
We also evaluate how each principal component (PC) of the reference dataset $X$ is captured as $Y$ is optimized for similarity. Specifically, for each PC $v_{i}$ of $X$, we use linear regression to find the optimal linear combination of columns of $Y$ that best predicts $v_{i}$. The goodness of fit is then quantified using the $R^{2}$ coefficient:
$R_{PC_{i}}^{2}:=1-\dfrac{\|(X-\hat{X})v_{i}\|^{2}}{\|Xv_{i}\|^{2}}$
where $\hat{X}=Y\hat{B}$ and $\hat{B}=\text{argmin}_{B}\|X-YB\|^{2}_{F}$.
All experiments were conducted on a NVIDIA GeForce GTX TITAN X and required less than a day of compute time.
B.3 Proof: theoretical analysis of CKA and NBS
Our results show that similarity in the high variance components seem to dominate the Centered Kernel Alignemnt (CKA) score. In contrast metrics like Normalized Bures Distance (NBS) seem to be more sensitive to changes in lower variance components. We explain this difference by showing, in theory, how CKA and NBS changes when perturbing a single principal component of the data. The result is that CKA depends quadratically on the variance of the perturbed principal component, whereas NBS has a linear dependence. We start the derivation from the matrix norm definitions of CKA and NBS.

### v3

@@ 4.2.2 Neural datasets
The optimization dynamics revealed in Figure 4a for a two-dimensional dataset also holds on real neural data (B.1). We now consider several neural datasets as the reference dataset for the optimization procedure. Figure 5a shows the optimization dynamics when a random noise Gaussian dataset is optimized towards the Siegel et al. 2015 dataset by maximizing angular Procrustes similarity (top) or CKA similarity (bottom). Each curve shows how the optimized dataset captures a single principal component of the reference dataset during the course of optimization; the yellow curve shows the highest variance component. Similar to the results in Figure 4, optimizing for angular Procrustes similarity, as opposed to CKA, captures more of the lower variance components in the data for a given similarity score.
@@
A convenient way of summarizing these curves is to note the similarity score required to capture a given principal component above some threshold. The PC threshold is defined here (and shown in Figure 5a for PC 1) as the centerpoint between the maximum and initial $R^{2}$ value for a given principal component. Figure 5b shows the score required to reach the principal component threshold for the different principal components in the Mante et al. 2013 dataset. Figures 5c and 5d show the same curves when the reference datasets are now the electrode recordings from Siegel et al. 2015 and Majaj et al. 2015. Figures 5b, 5c, and 5d additionally show that the metric version of CKA (angular CKA) increases CKA’s sensitivity to lower variance components.
@@
Neural curves predicted by matched Gaussians. Surprisingly, the optimization dynamics shown in these figures are well matched when the reference neural dataset is replaced by a new reference dataset that consists of random Gaussian numbers with variances for each principal component matched to the neural data (Figure 5e). The variance distribution of the reference dataset strongly determines when each principal component is captured during optimization.
4.3 Theoretical analysis of CKA and angular Procrustes
We aim to understand the increased sensitivity of CKA scores to high variance principal components compared to angular Procrustes. Following Harvey et al. (2023), we use the Normalized Bures Similarity (NBS) metric as an intermediary metric between CKA and angular Procrustes. Specifically, we study the sensitivity to dimensions of different variances by analyzing the similarity scores between a neural dataset and a modified version of the same dataset when a single principal component (PC) is perturbed (Figure 6). We “perturb” a single PC of the neural dataset by first decomposing the dataset into its projections onto every PC and then replacing a single one of these projections by Gaussian random noise that is rescaled so the variance is unchanged.
From NBS to CKA. NBS and CKA mainly differ by a choice of matrix norm as shown in the following formulation (Harvey et al., 2023) (see appendix C.1 for details):
@@ 4.3 Theoretical analysis of CKA and angular Procrustes
NBS quantifies similarity using the nuclear norm, which involves a sum of singular values, whereas CKA uses the Frobenious norm, which sums the square of the singular values. The additional square operation in CKA significantly increases the contribution of large variance components. We show this in theory by considering how CKA and NBS change when perturbing a single principal component of the data. The result is that CKA depends quadratically on the variance of the perturbed principal component, whereas NBS has a linear dependence (see proof in appendix C.2).
@@
where $\tilde{X}_{k}$ is equal to $X$ where only the $k^{\text{th}}$ principal component is modified, and $\lambda_{X}^{i}$ are the eigenvalues of $XX^{T}$.
We validate these results empirically by computing CKA and NBS scores between five different neural datasets and versions of these datasets when a single principal component is perturbed (see colored dots in Figure 6). As predicted by the theory, the NBS and CKA scores are well matched by, respectively, a linear and a quadratic function of the variance of the perturbed PC (see orange and blue curves).
From angular Procrustes to NBS. Harvey et al. (2023) showed that angular Procrustes is related to Normalized Bures Similarity (NBS) (Tang et al., 2020) by the arccosine function. We additionally normalize the angular Procrustes distance to obtain a score version, where 1 corresponds to perfect similarity.
[…]
The arccosine function decreases linearly around 00 with a slope of $-1$ and has a vertical asymptote in $1$. When NBS is small, which corresponds to high variances of the perturbed PC in Figure 6, angular Procrustes is also linear. When NBS is around 1, which corresponds to low variances of the perturbed PC in Figure 6, angular Procrustes has increased sensitivity compared to NBS, as explained by the large slope of arccosine around $1$.
Figure 6: The similarity score between five neural datasets and a modified version of these datasets when a single principal component is perturbed. The colored dots show the empirical CKA, NBS, and angular Procrustes scores. The solid lines show the predictions of our theory. In particular, the CKA scores decrease quadratically with the variance of the perturbed principal component, whereas NBS scores decrease linearly. Angular Procrustes is related to NBS by the arccosine function, which explains its linear dependence to high variance PCs and its increased sensitivity to low variance PCs compared to NBS.
Figure 7: We jointly optimized the values of both angular Procrustes and CKA or linear regression to illustrate the allowed ranges of both similarity scores (gray region enclosed by the solid black lines). (a) Different categories of possible relationships between a similarity measure A and a similarity measure B. (b) Allowable ranges of similarity scores for a fixed reference dataset, illustrated here for the Siegel 2015 dataset. These ranges are estimated by jointly optimizing a pair of similarity measures to get as close as possible to different target values (shown with the pink dots forming a square on the first plot). If angular Procrustes has a high score of 0.9 (horizontal dashed line) then linear regression will have a value above this. In contrast, a high linear regression score of 0.9 (vertical dashed line) does not imply a high angular Procrustes score, and a wide range of angular Procrustes scores are possible (see appendix C.3 for details).
@@ 4.4 Are metrics mutually independent?
A natural question that arises is whether these different similarity metrics are independent of each other, or if they exhibit consistent relationships. To address this, we consider three possible relationships between any two given similarity metrics (Figure 7). Independent: The two metrics are independent. A high score on one metric offers no guarantee of a high score on the other. Coupled: The two metrics are tightly coupled. A high score on one metric implies a high score on the other, and vice versa. Asymmetric: One metric subsumes the other. A high score on the first metric guarantees a high score on the second, but not the other way around.
We jointly optimize multiple similarity scores to find their allowed ranges (Figure 7) and show that a high angular Procrustes similarity implies a high CKA score, but not the converse. A high value of angular Procrustes implies a high score for unregularized linear regression but linear regression that is regularized and cross-validated across experimental conditions can take independent values.
@@ 5 Discussion
The central aim of our work is to explore the question of what it means for a similarity score to be good, as well as what drives a high similarity score. Our findings demonstrate that the interpretation of these scores is highly dependent on the specifics of both the metric and the dataset. We argue that similarity scores require further interpretation before making any claim based on them. Here we show two concrete ways of doing so, that incorporate the context of the dataset and similarity measure, by optimizing synthetic datasets to determine (i) how strongly task variables are encoded in order to reach the score (Figure 3), and (ii) how many PCs need to be captured in order to reach the score (Figures 4 and 5).
Limitations & future work. Our study focused on differentiable, geometry-based similarity measures and their optimization dynamics. Future work should investigate how it extends to other types of measures. Additionally, when we study the optimization dynamics we are studying both the behavior of the optimization method and the behavior of the similarity measure. Future work should investigate the potential limitations of our gradient-based optimization method, and how to better untangle it from the behavior of the similarity measure.
Moving forward, our python package aims to standardize similarity measures, facilitating a more cumulative scientific approach by centralizing findings related to these measures, and enabling more integrated benchmarking and comparisons. There are many similarity measures we have not analyzed in this work, for example, even CKA has at least 12 variations, which are easily accessible in our package. These similarity measures may, upon further investigation with the techniques we have proposed here, suffer from the same limitations we have demonstrated. Before using a similarity measure it is crucial to be aware of these limitations. Similarity package: https://github.com/nacloos/similarity-repository
6 Acknowledgments
We thank Laureline Logiaco for valuable discussions.
Appendix A Additional results: decode accuracy of task variables
Figures S1a and S1b show the decode accuracy of a linear classifier trained to decode task relevant variables for the Mante 2013 and the Siegel 2015 datasets (cross-validated across different conditions) as the similarity score increases. For both datasets, monkeys are shown a field of colored moving dots on each trial, and are required to attend to either color or motion information while ignoring the non-cued feature of the stimuli. Following Mante et al. (2013), we consider as the relevant task variables, the direction of motion of the dots, the color of the dots, the contextual cue, and the response of the monkey. We additionally binarize each of these variables for our decoding analysis. Before optimization, the synthetic datasets initially consisted of Gaussian noise and the decode accuracy was near the baseline chance level of 0.5 as expected for a binary classifier. We show results for the context task variable in Figure 3.
In Figure S2 we focus on a popular similarity measure, linear regression, and show how much task relevant information can be decoded from a synthetic dataset that is optimized to become increasingly more similar to the Siegel 2015 dataset. Even when linear regression is cross-validated and regularized, a high similarity score does not necessarily mean that task relevant information is encoded in a similar manner to neural data (leftmost panels).
Figure S1: (a, b) Decode accuracy for experimental variables versus similarity scores. Decode is from synthetic data optimized towards greater similarity with the neural data from (a) Mante et al. (2013) and (b) Siegel et al. (2015). Horizontal dashed lines indicate the decode accuracy from the neural data. (c, d) Decode accuracy from neural data versus number of principal components included in the decode. Decode uses data from (c) prefrontal cortex from Mante et al. 2013 and (d) FEF from Siegel et al. 2015.
Figure S2: Decode accuracy versus similarity scores when optimizing ridge regression scores with varying regularization strengths $\lambda$. Regression scores are cross-validated across conditions. Horizontal lines show the decode accuracy from the neural data for the task relevant variables, and colored curves shows the corresponding decode accuracy on a synthetic dataset that is optimized to become increasingly more similar to the neural data. For some values of the ridge regularization parameter, even at the highest similarity scores near 1, the synthetic dataset does not encode task relevant information to the same extent as the neural data.
Figure S3: We test whether removing low-variance principal components from the neural dataset increases the decoding accuracy, in order to see if this “denoising” may be a potential mechanism for explaining the increased decoding accuracy in the optimized dataset compared to the reference neural dataset in Figure 3. We show results for decoding of the context task variables from the Mante 2013 dataset after projecting onto principal components 1 through N, where principal component 1 is taken to have the highest variance, and N increases along the x-axis of the figure. The solid blue line shows the decoding accuracy from the synthetic dataset optimized with Angular Procrustes when the similarity score reaches a value of 0.9. Although the decoding accuracy from the neural data increases slightly when the lowest variance components are removed, it does not match the decoding accuracy from the synthetic data. This suggests the improved decoding accuracy from the synthetic data is probably a more complex form of denoising and would require further investigation.
Appendix B Neural Datasets and Models
B.1 Datasets
@@
  • Mante et al. (2013): Prefrontal cortex (PFC) electrode recordings during a contextual decision-making task involving colored moving dots. Activity is averaged across trials with the same condition (72 unique conditions) and averaged across 50 ms non-overlapping time bins during the 750 ms interval from 100 ms after dots onset to 100 ms after dots offset. The total number of neurons recorded is 727. To make optimization faster, we reduce the dimensionality of this dataset by keeping 99% of the variance, corresponding to the first 448 principal components. Data link111 https://www.ini.uzh.ch/en/research/groups/mante/data.html. The final dataset used in all the similarity analyses had size (time = 15, sample = 72, feature = 448).
  • Siegel et al. (2015): Electrode recordings from multiple cortical regions during a contextual decision-making task involving colored moving dots, similar to Mante et al. (2013). We analyze data from the Frontal Eye Field (FEF) region during the stimulus period of the task. Activity is averaged across trials with the same condition (42 unique conditions) and averaged across 50 ms non-overlapping time bins from 0 to 500 ms after dots onset. A total of 1220 neurons were recorded. We reduce the dimensionality of this dataset by keeping 99% of the variance, corresponding to the first 278 principal components. The final dataset used in all the similarity analyses had size (time = 11, sample = 42, feature = 278).
  • Hatsopoulos et al. (2007): Primary motor (M1) electrode recordings during a center-out reaching task. Simultaneous recordings were collected for 391 trials, 141 neurons. Activity is averaged across 50 ms non-overlapping time bins from 0 to 500 ms after movement onset. Data link [^2] https://datadryad.org/stash/dataset/doi:10.5061/dryad.xsj3tx9cm. The final dataset used in all the similarity analyses had size (time = 10, sample = 391, feature = 141).
  • Majaj et al. (2015): Inferior temporal (IT) electrode recordings during object image presentations. Simultaneous recordings for 3200 trials, 168 neurons. Activity is averaged across all time steps. The final dataset used in all the similarity analyses had size (sample = 3200, feature = 168).
  • Freeman et al. (2013): Primary (V1) and secondary (V2) visual area electrode recordings during texture and noise image presentations. Simultaneous recordings were collected for 135 trials, 205 neurons. Data is averaged across 150 ms non-overlapping time bins. The final dataset used in all the similarity analyses had size (time = 2, sample = 135, feature = 205).
We use the BrainScore [^3] https://github.com/brain-score/vision library (Schrimpf et al., 2020) for the Majaj et al. (2015); Freeman et al. (2013) datasets.
B.2 Recurrent Neural Network (RNN) models
RNN architectures. We show results for three commonly used RNN architectures: LSTMs (Hochreiter & Schmidhuber, 1997) and two choices for continuous time recurrent neural networks (CTRNNs), which differ by the position of the nonlinearity (Miller & Fumarola, 2012). A first alternative is given by the following equations.
@@ B.2 Recurrent Neural Network (RNN) models
We call it the LowPassCTRNN since it can be shown that its firing rate is a low-pass filter of the firing rate of the first architecture, which we call CTRNN (Miller & Fumarola, 2012). The RNNs are trained using supervised learning on simplified task inputs and outputs.
Comparing task-optimized RNNs to neural datasets. We train the RNNs on the tasks from Mante et al. (2013) and Siegel et al. (2015), and compare them to the respective neural datasets. We consider two neural datasets from prefrontal cortex (PFC) (Mante et al., 2013) and Frontal Eye Field (FEF) (Siegel et al., 2015) in monkeys performing an experimental task that required the animal to attend to either color or motion information while ignoring the non-cued feature of the stimuli. On each trial, a field of colored moving dots is shown. Monkeys are given a cue at the beginning of the trial to determine whether the dots in the stimulus are moving left vs right, or are red vs green. The monkey reported its choice with a saccade to one of two visual targets. In both datasets, we analyzed neural activity taken when the dot stimulus was presented. In panel a of Figure 2 neural activity is visualized in a low-dimensional space capturing task-relevant dynamics using the targeted dimensionality reduction method from Mante et al. 2013. Each curve shows the average neural activity for a different experimental condition. See Mante et al. 2013 for a detailed description of the analysis. This visualization highlights features of the data but the similarity scores were computed using the firing rates from the electrode recordings before any dimensionality reduction. In panel b of Figure 2 neural firing rates for two example neurons are shown with the colors denoting average.
@@
In Figure 2 c, d, we compare model-data similarity scores to two baseline scores. The Neuron-split baseline score is obtained by dividing the neurons from a single dataset into disjoint sets and then comparing. If the model-data scores are equal to the neuron-split scores this indicates that model activity is indistinguishable from the neural activity of other recorded neurons. The Condition-average baseline score is obtained by averaging the neural activity along the trial and condition dimensions and comparing it to the original dataset. Each neuron in this condition-averaged dataset still has a unique time-varying firing rate. This strong baseline shows the similarity to the original data that one can obtain by only keeping the condition-independent neural dynamics.
Appendix C Similarity Measures
C.1 Definitions
@@
Several methods have been proposed to measure similarity between models and neural data (see (Sucholutsky et al., 2023), (Klabunde et al., 2023) for comprehensive reviews). While some efforts have been made to characterize these metrics mathematically, for instance, by examining their invariance properties, clear guidance in interpreting similarity scores for a given scenario remains limited. To address this gap, we introduce a novel method for analyzing the specific aspects of data prioritized by different similarity metrics. Our method leverages the differentiability of these metrics and is applicable to a wide range of measures. We focus here on three commonly used metrics and their variants, linear regression (Yamins et al., 2014; Schrimpf et al., 2018), Centered Kernel Alignment (CKA) (Kornblith et al., 2019), and Procrustes distance (Williams et al., 2021; Ding et al., 2021). These methods quantify similarity based on the goodness of fit after aligning the representations. This alignment transformation allows for greater flexibility by making similarity measures invariant to specific transformations. For example, instead of demanding a one-to-one mapping of neurons between datasets, these methods can accommodate scenarios where neurons in one dataset correspond to linear combinations of neurons in the other.
@@
Unlike linear regression, CKA and Procrustes are symmetric, meaning the metric yields the same result whether applied to X, Y or Y, X. Moreover, they exhibit a different class of invariance. While linear regression is invariant under invertible linear transformations, CKA and Procrustes are invariant under orthogonal linear transformations. This stricter invariance to orthogonal transformations potentially reduces sensitivity to noise (Kornblith et al., 2019).
CKA measures the correlation between the kernels of two datasets $X$ and $Y$ (Kornblith et al., 2019). We consider here linear CKA where the kernels are $XX^{T}$ and $YY^{T}$.
@@
CKA scores range from 0 to 1, with 1 indicating perfect similarity. We also consider the arccosine of CKA, a metric satisfying the axioms of a distance metric (e.g., the triangle inequality) with 0 signifying perfect similarity (Williams et al., 2021). This metric, also known as Angular CKA (Lange et al., 2022), is then normalized to a similarity score between 0 and 1 for direct comparison with other scoring methods. This normalization involves dividing the angular distance by $\pi/2$ and subtracting the result from 1, ensuring the measure increases with similarity.
@@
Procrustes distance provides another approach for quantifying similarity (Williams et al., 2021; Ding et al., 2021). This metric identifies the optimal orthogonal alignment to maximize the correlation between $X$ and $Y$:
@@
where $\mathcal{O}$ is the group of orthogonal linear transformations. An angular distance metric can be obtained by taking the arccosine of this quantity (Williams et al., 2021).
As shown in Harvey et al. (2023), Procrustes distance is closely related to Normalized Bures Similarity (NBS) (Tang et al., 2020), with Procrustes distance equating the arccosine of NBS. NBS is defined as:
@@
This definition resembles CKA but utilizes the nuclear matrix norm instead of the Frobenius matrix norm. The Frobenius norm, denoted by $\|A\|_{F}$ for a matrix $A$, is calculated as the square root of the sum of squared singular values. The nuclear norm, $\|A\|_{*}$, is simply the sum of singular values. The implications of this difference in matrix norms for the weighting of principal components by CKA and NBS are further explored in C.2.
@@
C.2 Theoretical analysis of CKA and NBS
Our results show that similarity in the high variance components seem to dominate the Centered Kernel Alignment (CKA) score. In contrast, while the high variance components still have an overall larger impact than lower variance components, metrics like Normalized Bures Distance (NBS) seem to be more sensitive to changes in lower variance components relative to CKA. We explain this difference by showing, in theory, how CKA and NBS changes when perturbing a single principal component of the data. The result is that CKA depends quadratically on the variance of the perturbed principal component, whereas NBS has a linear dependence. We start the derivation from the matrix norm definitions of CKA and NBS (Harvey et al., 2023).